OpenStudy (anonymous):

f(x)=square root 4x+2, g(x)=square root 25x-4 Find (fg)(x)

OpenStudy (anonymous):

is that f times g or f of g?

OpenStudy (anonymous):

\[fg(x)=\sqrt{4x+2}\times \sqrt{25x-4}=\sqrt{(4x+2)(25x-4)}\]

OpenStudy (anonymous):

\[f(x)=\sqrt{4x+2}, g(x)=\sqrt{25x-4}\]

OpenStudy (anonymous):

f times g